2(3x-4)=3+6x
You would use the distributive property for the left part of the problem.
1. 2*3x 2*-4
2. 6x - 8 = 3 + 6x
3. 6x - 8 -3 = 3 -3 + 6x
4. 6x - 11 = 6x
5. 6x - 6x - 11 = 6x - 6x
6. -11 = 0

Found 2 solutions by umail08, vtskge:
Answer by umail08(87)

Please check to see if you wrote the problem correct.

Answer by vtskge(1)

2(3x-4)=3+6x
You would use the distributive property for the left part of the problem.
1. 2*3x 2*-4
2. 6x - 8 = 3 + 6x
3. 6x - 8 -3 = 3 -3 + 6x
4. 6x - 11 = 6x
5. 6x - 6x - 11 = 6x - 6x
6. -11 = 0